1) Assume that for olive oil in Palestine, the price elasticity of demand is 0.8 and the price



elasticity of supply is 0.7. And suppose that the demand for the Palestinian olive oil has



increased by 20%. Do you think the price of Palestinian olive oil would increase or decrease?



By how much? Explain and show calculations.

1
Expert's answer
2021-12-06T09:37:03-0500

Price elasticity of demand (PED) = 0.8

% change in quantity demanded = 20%


"PED=\frac{\bigtriangleup Q}{\bigtriangleup P}"

"0.8=\frac{0.20}{\bigtriangleup P}"


"\bigtriangleup P=\frac{0.20}{0.8}=0.25"


This means the price of Palestinian olive oil would decrease by 25% that resulted as increase in quantity demanded by 20%.

Based on this, it can say that the price elasticity of demand is inelastic for Palestinian olive oil.